# Omni-ID: Holistic Identity Representation Designed for Generative Tasks

### v1

@@ Abstract
We introduce Omni-ID, a novel facial representation designed specifically for generative tasks. Omni-ID encodes holistic information about an individual’s appearance across diverse expressions and poses within a fixed-size representation. It consolidates information from a varied number of unstructured input images into a structured representation, where each entry represents certain global or local identity features. Our approach uses a few-to-many identity reconstruction training paradigm, where a limited set of input images is used to reconstruct multiple target images of the same individual in various poses and expressions. A multi-decoder framework is further employed to leverage the complementary strengths of diverse decoders during training. Unlike conventional representations, such as CLIP and ArcFace, which are typically learned through discriminative or contrastive objectives, Omni-ID is optimized with a generative objective, resulting in a more comprehensive and nuanced identity capture for generative tasks. Trained on our MFHQ dataset – a multi-view facial image collection, Omni-ID demonstrates substantial improvements over conventional representations across various generative tasks.
@@ 1 Introduction
Furthermore, existing face representation methods, typically derived from networks optimized for discriminative tasks such as ArcFace [7] or text-aligned image encoders like CLIP [21], are not well suited for generative applications. While these models effectively distinguish between individuals or extract semantic features, they are constrained by a discriminative bottleneck that omits fine-grained identity details. Consequently, they struggle to capture the subtle nuances that define a person’s unique identity, especially across different poses and expressions, as demonstrated in Fig. 2.
@@
Our Omni-ID encoder is transformer-based and employs a fixed-size set of learnable queries to produce a consistent, fixed-size representation of an individual’s identity. This design ensures that regardless of the number of input images provided, the encoder yields a structured encoding by blending the input images via keys and values throughout attention layers. The fixed-size representation is essential for downstream applications, as it establishes a ‘structured’ encoding where each feature within the representation can correspond to specific identity attributes, such as different facial regions as visualized in Fig. 10. Structured representations allow downstream tasks to focus on learning from the distilled identity features without being distracted by the noise and variability present in individual input images.
To validate Omni-ID’s effectiveness, we conduct extensive experiments comparing our method against state-of-the-art baselines, including ArcFace [7] and CLIP [21] representations. Notably, with our representation the generation quality scales with the number of input images, enabling it to capture a more comprehensive view of the individual, as shown in  Fig. 2. In addition, we demonstrate Omni-ID’s superiority in two widely used face generative tasks: controllable face synthesis and personalized text-to-image generation. Extensive experiments show that Omni-ID significantly improves identity preservation across a range of poses, expressions and contexts, achieving higher fidelity in generating photorealistic images.
Figure 3: Gallery of Omni-ID in personalized T2I generation. Omni-ID enables high identity preservation. Results achieved by injecting Omni-ID representation through IP-Adapter [37] into the frozen FLUX dev model [3] without LoRA [16].
@@ 2 Related work
Face representation provides a foundation for 3D face reconstruction, accurately distinguishing identities and synthesizing realistic face images. Parametric 3D Morphable Models [4, 5, 18] have been historically used to represent face shape geometry through identity and expression blendshapes combined with pose. However, these representations are coarse and lack the appearance details needed for photo-realistic generation. In recognition tasks, approaches like CosFace [32] and ArcFace [7] have improved identity discrimination by utilizing a margin loss, which enhances intra-class compactness and inter-class separability, and have also been widely applied in generation tasks. More recently, FaRL [40] creates a more descriptive facial token representations by fine-tuning the pretrained visual model, CLIP [21] on large-scale face-text paired datasets. In contrast to these discriminative or contrastive facial identity representations, our Omni-ID is optimized with a generative objective, resulting in a more nuanced identity representation well suited for generative tasks.
Face synthesis has evolved significantly, starting with StyleGAN [17], which set new standards for high-quality, realistic face generation through a well-structured latent space. However, StyleGAN offered limited control over individual facial features. To address this, researchers introduced methods to invert real images into StyleGAN’s latent space [1, 13, 2, 30, 22], enabling attribute manipulation by altering latent codes. Given recent advances in generative models, diffusion [15, 29, 8, 14] and flow-based [23] generative models offer even higher quality generations than GAN-based approaches. In Sec. 4.2 we study controllable face generation and show improved identity fidelity when using our proposed face representation.
@@ 3.2 Few-to-Many Identity Reconstruction
where $\tilde{x}^{r}_{i}$ is the corrupted target image. Intuitively, a corruption process destroys information from the target image, and prevents the decoder from utilizing the target image to achieve autoencoding. This forces the decoder to rely on the encoder feature $\ell$ to infer identity information, thereby encouraging the encoder to learn a robust identity representation. Meanwhile, the corrupted target image still retains cues about other conditions, such as lighting, pose, and subtle hints of expression, providing essential context that, when combined with representation from the encoder, enables accurate reconstruction. To leverage the strengths of different corruption types, we employ distinct objectives, which we detail below.
@@ 4 Experiments
The first task is controllable face generation (Sec. 4.2), where a downstream generator produces an image of an individual in unseen poses based on an identity representation and a target pose (i.e., landmarks). This task tests the representation’s ability to capture nuanced changes with varying poses and expressions.
The second task is personalized text-to-image generation (Sec. 4.3). Here, the generator creates scene-level images that maintain both individual identity and the quality of the original text-to-image model. Our results on this task further validate that the Omni-ID representation outperforms existing alternatives and demonstrates its effectiveness in the popular application of personalized generation.
Lastly, we validate our design choices including each of the few-to-many identity reconstruction training paradigm, the decoding objectives, our proposed dataset, and their hyperparameters (Sec. 4.4).
@@ 4.1 Implementation Details
Table 1: Quantitative comparisons to different representations on controllable face generation. The backbone is the same for all methods: IP-Adapter + ControlNet. All baselines undergoes Flow-Matching pretraining to initialize IP-Adapters to converge for fair comparison. We show three results of using $1/3/5$ inputs.
[…]
ArcFace 0.515 / 0.523 / 0.529 2.4 / 2.3 / 2.3
CLIP 0.648 / 0.670 / 0.680 2.3 / 2.3 / 2.2
ArcFace + CLIP 0.638 / 0.655 / 0.663 2.4 / 2.4 / 2.3
Omni-ID (Ours) 0.708 / 0.728 / 0.737 2.1 / 2.0 / 2.0
@@
Omni-ID training. Our MTD decoder is trained for $250$K steps with a masking ratio of $95\%$. Our Flow-Matching Decoder is trained for $10$K steps using FLUX dev [3] as the base model. Both stages are trained in MFHQ with $44$ held out videos as testing and others as training. See Appendix for further training details.
Baseline representations. In the controllable face generation task, we compare our Omni-ID representation to other commonly used identity representations: pretrained CLIP features [21] and Arc-Face embedding [7]. We also compare to CLIP+ArcFace, following FaceIDPlus [37], where the ArcFace embeddings are projected into queries and CLIP features are used as keys and values to get the representation through the same attention mechanism as our Omni-ID. For CLIP representations, we use all $257$ tokens for all baselines. For ArcFace, we project embedding $\mathbb{R}^{1\times 512}$ to $256$ tokens $\mathbb{R}^{256\times 512}$ for better quality. For the personalized text-to-image gneeration task, we compare our Omni-ID representation to pretrained CLIP embedding. For both tasks, we ensure fair comparisons with the baselines by training the downstream models to convergence when using each of the representations (detailed in Appendix).
Metrics. We use ID similarity and pose error as the evaluation metrics in controllable face generation. ID similarity is measured using the standard cosine similarity between the ArcFace features extracted from the generation and ground truth. Pose error [25] is measured by the sum of absolute differences of yaw, pitch, roll in degrees.
Figure 8: Qualitative comparisons with different representations in personalized T2I generation. We show results of the same IP-Adapter trained with different representations. Our Omni-ID achieves better ID preservation for both single and multiple input images. See more examples and how IP-Adapter+Omni-ID significantly outperforms other personalization methods [34, 19] in Appendix.
@@ 4.2 Controllable Face Generation
Given a pretrained representation, we train a combination of ControlNet [39] and IP-Adapter with frozen FLUX for controllable face generation. The ControlNet receives as input landmark images of the target pose, where IP-Adapter injects frozen face representations. Methods are evaluated on the test set of MFHQ with 44 identities from the held-out videos. Compared to the ArcFace, CLIP, or their combined representations, the Omni-ID representation shows better performance in both identity preservation and pose accuracy (see Tab. 1). Beyond metrics, Fig. 7 highlights qualitative differences between representations using 5 inputs driven by template landmarks. While ArcFace encodes facial features effectively for recognition tasks, it is overly invariant to attributes like age and skin tone. CLIP preserves general visual features but struggles with adaptivity to new poses and expressions due to its instance-level encoding and lack of fine-tuning for facial details. Consequently, facial features such as beards (last row) are not accurately represented in CLIP, and sensitivity to pose and expression changes is noticeable. In contrast, Omni-ID achieves high-fidelity identity preservation, capturing detailed facial characteristics across diverse pose and expressions.
@@ 4.3 Personalized T2I Generation
Given a pretrained representation, we train an IP-Adapter to inject into frozen FLUX-dev. Notice, this differs from the denoising decoder we used during representation learning in its data. Here, the input data is an image of the face, but the target image is a scene-level image and has a corresponding text caption. We train all baselines in the same internal licensed image dataset ($\sim 1$M single-view images), and evaluate them on $10$ identities and $20$ diverse prompts. ID similarity is employed as the metric. See Appendix for the quantitative results. As can be seen in Fig. 8, Omni-ID representation demonstrates superior performance in terms of identity preservation when applied to personalized text generation, outperforming CLIP in both single input image, as well as in the the multiple input images case. See Fig. 3 for more results of Omni-ID. See Appendix for more qualitative comparisons using different base models.
@@ 4.4 Ablations & Analyses
Tab. 2 summarizes experimental results where we used face generation qualities from our Flow-Matching decoder as evaluation for different MTD pre-training configurations. We ablated the number of images used as inputs and targets during MTD pre-training, removing MTD completely and changing other hyperparameters. Tab. 3 summarizes experimental results of the downstream controllable face generation performance by using different checkpoints obtained by ablating the individual decoding objectives and using an alternative dataset [42] instead of our MFHQ.
@@ 5 Conclusions
We introduced Omni-ID, a facial identity representation tailored for generative tasks, which captures an individual’s holistic appearance across various expressions and poses. Trained in a few-to-many identity reconstruction framework, the Omni-ID encoder encodes fine-grained facial features from diverse input images, demonstrating superior identity preservation. Unlike discriminative representations like ArcFace and CLIP, Omni-ID retains nuanced identity information critical for high-fidelity generative tasks.
@@
Our results suggest that generative task-based identity representation holds transformative potential for diverse facial generation applications. We anticipate that this approach will inspire further innovation, broadening the capabilities and scope of generative identity modeling across a wider range of applications. Improvements in dataset scale and consistency, as well as the number and type of the decoders would further enhance robustness. Additionally, Omni-ID does not represent attributes that are not intrinsic to the face, such as hair, which can result in these features being “hallucinated” in downstream tasks. Extending Omni-ID to include a more comprehensive set of attributes remains an open direction for future research.
@@
Acknowledgement. The authors would like to acknowledge other members of the Snap Creative Vision team for valuable feedback and insightful discussions throughout this project.

### v2

@@ Abstract
We introduce Omni-ID, a novel facial representation designed specifically for generative tasks. Omni-ID encodes holistic information about an individual’s appearance across diverse expressions and poses within a fixed-size representation. It consolidates information from a varied number of unstructured input images into a structured representation, where each entry represents certain global or local identity features. Our approach uses a few-to-many identity reconstruction training paradigm, where a limited set of input images is used to reconstruct multiple target images of the same individual in various poses and expressions. A multi-decoder framework is introduced to leverage the complementary strengths of diverse decoders during training. Unlike conventional representations, such as ArcFace and CLIP, which are typically learned through discriminative or contrastive objectives, Omni-ID is optimized with a generative objective, resulting in a more comprehensive and nuanced identity capture for generative tasks. Trained on our MFHQ dataset – a multi-view facial image collection, Omni-ID demonstrates substantial improvements over conventional representations across various generative tasks.
@@ 1 Introduction
Furthermore, existing face representation methods, typically derived from networks optimized for discriminative tasks such as ArcFace [10] or text-aligned image encoders like CLIP [29], are not well suited for generative applications. Intuitively, Information Bottleneck Principle [3] suggests that subtle variations critical for generation but irrelevant to class boundary are likely lost in discriminative training. Consequently, discriminative features struggled to capture the subtle nuances that define a person’s unique identity especially across different poses and expressions. Refer to Figs. 2, 7 and 6 for examples of an inaccurate nose, missing beards, and an inaccurate head shape when using discriminative features for generation.
@@
Our Omni-ID encoder is transformer-based and employs a fixed-size set of learnable queries to produce a consistent, fixed-size representation of an individual’s identity. The fixed-size representation is essential for downstream applications, as it establishes a ‘structured’ encoding where each feature within the representation can correspond to specific identity attributes, such as different facial regions as visualized in Fig. 10. Structured representations allow downstream tasks to focus on learning from the distilled identity features without being distracted by the noise and variability present in individual input images.
To validate Omni-ID’s effectiveness, we conduct extensive experiments comparing our method against state-of-the-art baselines, including ArcFace [10] and CLIP [29] representations. Notably, with our representation the generation quality scales with the number of input images, enabling it to capture a more comprehensive view of the individual, as shown in Fig. 2. In addition, we demonstrate Omni-ID’s superiority in two widely used face generative tasks: controllable face synthesis and personalized text-to-image generation. Extensive experiments show that Omni-ID significantly improves identity preservation across a range of poses, expressions, and contexts, achieving higher fidelity in generating photorealistic images.
@@ 2 Related work
Face representation provides a foundation for 3D face reconstruction, accurately distinguishing identities and synthesizing realistic face images. Parametric 3D Morphable Models [5, 6, 24] have been historically used to represent face shape geometry through identity and expression blendshapes combined with pose. However, these representations are coarse and lack the appearance details needed for photo-realistic generation [13, 9, 30]. In recognition tasks, approaches like CosFace [41] and ArcFace [10] have improved identity discrimination by utilizing a margin loss, which enhances intra-class compactness and inter-class separability, and have also been widely applied in generation tasks. More recently, FaRL [50] creates a more descriptive facial token representations by fine-tuning the pretrained visual model, CLIP [29] on large-scale face-text paired datasets. In contrast to these discriminative or contrastive facial identity representations, our Omni-ID is optimized with a generative objective, resulting in a more nuanced identity representation well suited for generative tasks.
Face synthesis has evolved significantly, starting with StyleGAN [22], which set new standards for high-quality, realistic face generation through a well-structured latent space. However, StyleGAN offered limited control over individual facial features. To address this, researchers introduced methods to invert real images into StyleGAN’s latent space [1, 18, 2, 39, 31], enabling attribute manipulation by altering latent codes. Given recent advances in generative models, diffusion [20, 38, 11, 19, 26] and flow-based [32] generative models offer even higher quality generations than GAN-based approaches. In Sec. 4.2 we study controllable face generation and show improved identity fidelity when using our proposed face representation.
@@ 3.2 Few-to-Many Identity Reconstruction
where $\tilde{x}^{r}_{i}$ is the corrupted target image. Intuitively, a corruption process destroys information from the target image, and prevents the decoder from utilizing the target image to achieve autoencoding. This forces the decoder to rely on the encoder feature $\ell$ to infer identity information, thereby encouraging the encoder to learn a robust identity representation. Meanwhile, the corrupted target image still retains cues about other conditions, such as lighting, pose, and subtle hints of expression, providing essential context that, when combined with representation from the encoder, enables accurate reconstruction. To leverage the strengths of different corruption types, we employ distinct objectives, all with few-to-many identity reconstruction, which we detail below.
@@ 3.3 Multi-Decoder Objectives
Table 1: Quantitative comparisons to different representations on controllable face generation. The backbone is the same for all methods: IP-Adapter + ControlNet. All baselines undergoes Flow-Matching pretraining to initialize IP-Adapters to converge for fair comparison. We show results of using different number of inputs in two test sets.
Dataset MFHQ Test Webface Test [27]
[…]
#inputs 1 / 3 / 5 / 7 1 / 3 / 5 / 7 3 / 5 / 8 / 16 3/ 5 / 8 / 16
ArcFace 0.515 / 0.523 / 0.529 / 0.535 2.4/2.3/2.3/2.3 0.379 / 0.373 / 0.370 / 0.371 / 0.373 2.5/2.4/2.5/2.4/2.4
CLIP 0.648 / 0.670 / 0.680 / 0.682 2.3/2.3/2.2/2.2 0.695 / 0.696 / 0.696 / 0.695 / 0.695 2.9/2.8/2.8/2.7/2.8
ArcFace + CLIP 0.638 / 0.655 / 0.663 / 0.664 2.4/2.4/2.3/2.3 0.652 / 0.654 / 0.656 / 0.658 / 0.655 2.9/2.8/2.8/2.7/2.8
Omni-ID (Ours) 0.708 / 0.728 / 0.737 / 0.742 2.1/2.0/2.0/2.0 0.774 / 0.779 / 0.781 / 0.784 / 0.785 2.7/2.6/2.6/2.6/2.6
@@ 4 Experiments
The first task is controllable face generation (Sec. 4.2), where a downstream generator produces an image of an individual in unseen poses based on an identity representation and a target pose (i.e., landmarks). This task tests the representation’s ability to capture nuanced changes with varying poses and expressions. The second task is personalized text-to-image generation (Sec. 4.3). Here, the generator creates scene-level images that maintain both individual identity and the quality of the original text-to-image model. Lastly, we validate our design choices including each of the few-to-many identity reconstruction training paradigm, the decoding objectives, our proposed dataset, and their hyperparameters (Sec. 4.4).
@@ 4.1 Implementation Details
Omni-ID training. Our MTD decoder is trained for $200$K steps with a masking ratio of $95\%$. Our Flow-Matching Decoder is trained for $10$K steps using FLUX dev [4] as the base model. Both stages are trained in MFHQ with $44$ held out videos as testing and others as training.
Baseline representations. In the controllable face generation task, we compare our Omni-ID representation to other commonly used identity representations: pretrained CLIP features [29] and ArcFace embedding [10]. We also compare to CLIP+ArcFace, following FaceIDPlus [47], where the ArcFace embeddings are projected into queries and CLIP features are used as keys and values to get the representation through the same attention mechanism as our Omni-ID. For CLIP representations, we use all $257$ tokens for all baselines. For ArcFace, we project embedding $\mathbb{R}^{1\times 512}$ to $256$ tokens $\mathbb{R}^{256\times 512}$ for better quality. For both tasks, we ensure fair comparisons with the baselines (see Appendix).
Metrics. We use ID similarity as metric, which is measured using the cosine similarity between the face recognition features [12] extracted from the generation and ground truth. In controllable face generation, pose error [34] metric is also provided which is measured by the sum of absolute differences of yaw, pitch, roll in degrees.
Figure 7: Qualitative comparisons with different representations in personalized T2I generation. We show results of the same IP-Adapter trained with different representations. Omni-ID achieves better ID preservation for both single and multiple input images. See more examples and how Omni-ID significantly outperforms other personalization methods [43, 25] in Appendix.
@@ 4.2 Controllable Face Generation
Given a pretrained representation, we train a combination of ControlNet [49] and IP-Adapter with frozen FLUX for controllable face generation. ControlNet receives as input target landmark pose, where IP-Adapter injects frozen face representations. The methods are evaluated in two test sets: (1) all identities from Webface21M [52, 27] consisting of at least 16 photos with minimum ID similarity 0.6 and minimum pose differences $7\degree$, and (2) MFHQ test set. Compared to ArcFace, CLIP, or their combined representations, Omni-ID shows better performance in identity preservation in  Tab. 1. All methods achieve similar pose errors indicating their convergence. Beyond metrics, Fig. 6 highlights qualitative differences between representations using 5 inputs driven by template landmarks. Although ArcFace encodes facial features effectively for recognition tasks, it is overly invariant to attributes such as age and skin tone. CLIP preserves general visual features but struggles with adaptivity to new poses and expressions due to its instance-level encoding and lack of fine-tuning for facial details. Consequently, facial features such as beards (last row) are not accurately represented in CLIP, and sensitivity to pose and expression changes is noticeable. In contrast, Omni-ID achieves high-fidelity identity preservation, capturing facial details across diverse poses and expressions.
Figure 8: Gallery of Omni-ID in personalized T2I generation. Omni-ID enables high identity preservation. Results achieved by injecting Omni-ID representation through IP-Adapter [47] into the frozen FLUX dev model [4] without LoRA [21] or postprocessing.
@@ 4.3 Personalized T2I Generation
Given a pretrained representation, we train an IP-Adapter to inject into frozen FLUX-dev. Notice, this differs from the denoising decoder we used during representation learning in its data. Here, the input data is an image of the face, but the target image is a scene-level image and has a corresponding text caption. We train all baselines in the same internal licensed image dataset ($\sim$1M single-view images), and evaluate them on $10$ identities and $20$ diverse prompts. ID similarity is employed as the metric. See Appendix for the quantitative results. As can be seen in Fig. 7, Omni-ID representation demonstrates superior performance in terms of identity preservation when applied to personalized text generation, outperforming CLIP in both single input image, as well as in the the multiple input images case. See Fig. 8 for more results of Omni-ID. See Appendix for more qualitative comparisons using different base models e.g. SD [33].
@@ 4.4 Ablations & Analyses
Tab. 2 ablates MTD pre-training configurations through face generation qualities by the Flow-Matching decoder. Tab. 3 ablates pre-training objective and dataset in the task of downstream controllable face generation. We present results using 1 or 3 input images in the ablation study, as additional inputs yield only marginal performance improvements as aforementioned and the models are trained with at most 3 inputs.
@@ 5 Conclusions
We introduced Omni-ID, a facial representation tailored for generative tasks, which captures an individual’s holistic appearance across various expressions and poses. Trained in a few-to-many identity reconstruction framework with a multi-decoder objective, Omni-ID encodes a fixed-size tokenized representation from diverse unstructured input images, demonstrating superior identity preservation. Unlike discriminative representations like ArcFace and CLIP, Omni-ID retains nuanced identity information critical for high-fidelity generation. Moreover, the quality of Omni-ID improves with an increasing number of input images.
@@
Our results suggest that generative identity representation holds transformative potential for diverse facial generation applications. We anticipate that our approach will inspire further innovation, broadening the capabilities and scope of generative identity modeling across a wider range of applications. Improvements in dataset scale and consistency, as well as the number and type of the decoders would further enhance robustness. Collection of large-scale multiple photos of the same identity in a more diverse context or lighting augmentation will also improve the robustness of Omni-ID, addressing lighting injection and skin tone predominance issues. Additionally, Omni-ID does not represent attributes that are not intrinsic to the face, such as hair, which can result in these features being “hallucinated” in downstream tasks. Extending Omni-ID to include a more comprehensive set of attributes remains an open direction.
@@
Acknowledgement. The authors would like to acknowledge Jian Wang, Qiang Gao, and Sizhuo Ma from the Computational Imaging team at Snap for their advices with face quality assessment. We also thank other members of the Snap Creative Vision team for their valuable feedback and insightful discussions throughout the course of this project.
Appendix A Experiment Details
A.1 Baseline Details
CLIP. Throughout our experiments, we use CLIP-H from OpenAI [29] as the feature extractor, which works slightly better than CLIP-B/L. We use the full representation, i.e. $257$ tokens ($256$ spatial tokens with $1$ class token) from the second last layer following IP-Adapter Full [47], which improves ID preservation compared to using only class token or a reduced number of tokens (e.g. $16$). For multiple inputs, we use token concatenation following IP-Adapter, which outperforms simple averaging. For fair comparisons, we train IP-Adapter with CLIP representation in the same Flow Matching Decoder stage for 5K steps with an effective batch size $32$ (roughly $1$ epoch in MFHQ). The convergence happens at around $4$K steps.
ArcFace. We use the ArcFace [10] model from insightface [8] throughout the experiments. We project Arcface embedding from $\mathbb{R}^{1\times 512}$ to 256 tokens with $1280$ channels $\mathbb{R}^{256\times 1280}$, which is comparable to CLIP and Omni-ID in terms of representation size. Using $256$ tokens improves its ID preservation compared to using $4$ or $16$ tokens only in IP-Adapter FaceID [47], and outperforms other reduced number of tokens ($64$). We concatenate representations in the token dimension for multiple inputs, where averaging merging also reaches a similar results for ArcFace representation. We train IP-Adapter with ArcFace representation in the Flow Matching Decoder stage by $75$K steps to converge. Compared to CLIP and Omni-ID which take about $5$K steps, the convergence of ArcFace is rather slow, due to its over-compactness for generative tasks.
ArcFace+CLIP. Following IP-Adapter FaceIDPlus [47], Arc-Face+CLIP baseline projects ArcFace tokens from the average ArcFace embeddings in $\mathbb{R}^{1\times 512}$ to $256$ queries in $\mathbb{R}^{256\times 1280}$, where each individual CLIP features in $\mathbb{R}^{257\times 1280}$ are used as keys and values to aggregate multiple inputs. For a fair comparison to Omni-ID, the same transformer with self attention layers is used to merge features. Both the improved number of tokens and the self-attention layers in transformer improves the face quality compared to the original implementation in FaceIDPlus, where $4$ or $16$ tokens are used as query and Q-former [23] without self attentions are employed.
A.2 Decoders and Training Details
Decoders details.
  • Masked transformer decoder. MTD is built by $6$ CA blocks and $2$ SA blocks, which reaches high-quality reconstruction while smaller number of decoder blocks might compensate encoder quality due to the lower decoding ability. Mask ratio of MTD is set to $95\%$, i.e. 5% patches are visible during training, which leads better encoder performance in downstream tasks than mask ratio $85\%$ or $99\%$. The patch size for the decoder is set to $14\!\times\!14$ to balance the speed and quality.
  • Flow-matching denoising decoder. For the Flow-Matching Decoder, FLUX dev [4] serves as the base model. We implement a FLUX-based version of IP-Adapter [47], where the Omni-ID representation is injected into all blocks, including MM-DiT and DiT blocks, via learnable decoupled attention layers. Injecting into both block types results in slightly better quality compared to injecting into only MM-DiT blocks or only DiT blocks, although this improvement is not critical. Each decoupled attention layer optimizes a single linear projection to map $\ell$ from $\mathbb{R}^{L\times C}$ to $\mathbb{R}^{L\times 3250}$, where $3250$ is the channel size used in FLUX. During the Flow-Matching Decoder stage, the Omni-ID encoder and the projection layers of the decoupled attention layers are optimized, while the original parameters in FLUX remain frozen.
Training Details. Omni-ID uses a two-stage few-to-many identity reconstruction training process: the MTD stage and the Flow Matching Decoder stage. The MTD stage is trained on our MFHQ dataset at an image resolution of $448$ using a constant learning rate of $1e^{-4}$, an effective batch size of $256$ (distributed as $32$ batches across $8$ NVIDIA A100 GPUs), and the AdamW optimizer for 250K iterations. The Flow Matching Decoder stage is trained on the same dataset at a resolution of $512$, with a constant learning rate of $1e^{-4}$, an effective batch size of $32$, and the AdamW optimizer for 5K iterations. In both stages, we uniformly sample a variable number of inputs ($1$ to $3$) and generate all $8$ targets for each identity.
Downstream Details.
  • Controllable face generation. For all experiments, we freeze the face representation encoder and optimize both the ControlNet and IP-Adapter using a constant learning rate of $2e^{-5}$ and an effective batch size of $16$ for $15$K steps. The models are trained on MFHQ with a variable number of inputs (uniformly sampled between $1$ and $7$) and a single target image, all at a resolution of $512\times 512$. All models converge well before reaching $15$K steps. The ControlNet is implemented and initialized as described in [46]. The IP-Adapter is initialized from our Flow Matching Decoder. For fair comparisons, other representations (e.g., CLIP and ArcFace) also undergo the Flow Matching Decoder training stage to achieve convergence, requiring $5$K steps for CLIP and $75$K steps for ArcFace. In the benchmark, ground truth landmarks from the same identity are used as ControlNet inputs, and metrics are calculated between the generated images and the targets. The generation resolution is set to $512\times 512$.
  • Personalized T2I generation. We integrate frozen face representations into the frozen FLUX dev base model [4] using learnable decoupled attentions, following the approach outlined in IP-Adapter [47]. Injecting into MM-DiT blocks is unnecessary in personalized T2I and does not affect the image quality. The IP-Adapter is trained using a simple flow-matching loss without additional regularization (e.g. ID loss, alignment loss [17]) and without employing LoRA [21]. These regularization and LoRA modules are left for future study as orthogonal to our work. Our training is performed at a resolution of $512\times 512$ for 50K steps with a constant learning rate of $1e^{-4}$, using the AdamW optimizer. Subsequently, we fine-tune the IP-Adapter at a resolution of $768\times 768$ for 20K steps, maintaining the same hyperparameters. Models are trained on our internal purchased dataset (Getty Images). For fair comparisons, other representations, such as CLIP and ArcFace, are trained under the same hyperparameters unless otherwise noted. Due to its slower convergence compared to Omni-ID and CLIP, ArcFace requires 100K steps in the first stage to achieve convergence. Inference for this task is performed at a resolution of $1024\times 1024$.
MFHQ Details. Refer to Fig. I how MFHQ is collected for each video clip.
Figure I: Illustration of MFHQ Creation. Given a video, we first detect faces and distribute identities to different clips by a threshold based on the cosine distance of face embeddings [10]. Then, a face quality estimation [7] is applied to sort the quality of frames within each identity. 20% faces with lowest quality are removed. A head pose estimation [34] is employed to estimate the poses for each face which are used to cluster the frames into $M=16$ clusters. Finally, $8$ frames are sampled $M$ clusters, where each cluster is only sampled at most once. The sum of absolute pose differences is assured larger than $15$ degree for each pair.
Appendix B Supplementary Experiments
B.1 Additional Controllable Face Generation
Fig. II further compares Omni-ID with ArcFace [10] and CLIP [29] in the context of controllable face generation. Unlike the benchmark case presented in the main paper, where Ground Truth landmarks were used to guide identity-specific generation, here we use the template-driven landmarks as conditions. 9 template images are collected to obtain a grid of expression and pose in FLAME code [24] through 3D mesh reconstruction by 3D landmark estimation. Then, we use the FLAME shape code for each identity with the template FLAME expression and pose code from each template to get the rigged mesh. From each mesh, 2D lanmarks are rendered as condition to generate each view at the grid for each identity.
While CLIP demonstrates strong baseline performance, it struggles with identity preservation and fails to generate realistic faces when the pose and expression differ significantly from the input images. This limitation arises because CLIP is an instance-level representation model. In contrast, our Omni-ID is an identity-level representation, specifically trained to reconstruct faces in new poses and expressions. Consequently, Omni-ID achieves significantly better identity preservation while generating new faces of the identity.
B.2 Additional Personalized Text-to-Image
Table I: Quantitative comparisons to the state-of-the-art on personalized T2I generation. ID Similarity are computed by the cosine distance between the generated samples and the five images of each identity. We compute the average and std across identities. The base models are FLUX [4] for all methods.
[…]
IPA-FaceID 0.3535$\pm$0.1982
IPA-FaceIDPlusV2 0.4327$\pm$0.1090
IPA-Full 0.6649$\pm$0.0846
PuLID 0.7289$\pm$0.0572
IPA-Omni-ID Schnell (Ours) 0.7306$\pm$0.0793
IPA-Omni-ID (Ours) 0.8026$\pm$0.0421
Compare to State-of-the-Art. We compare Omni-ID+IP-Adapter (IPA Omni-ID) to the state-of-the-art IP-Adapter [47], InstantID [43], PhotoMakerV2 [25], PuLID [17] in Fig. III and Fig. IV when using FLUX Dev [46] and Stable Diffusion (SD) [33] as the base model, respectively. Our IPA Omni-ID trained by the simple flow matching loss without any advanced techniques such as LoRA [21], ID loss [17], aligment loss [17], stacked embedding [25], IdentityNet [43], achieves the highest ID preservation. Refer to gallery.m4v for all visual results of our model. Tab. I compares IPA Omni-ID with the state-of-the-art personalized T2I employed FLUX as the base model. Our IPA Omni-ID outperforms others with the highest identity similarity.
Beyond FLUX Dev Experiments. Despite Omni-ID is trained using FLUX dev [4] as the Flow Matching Decoder, Omni-ID can be applied to any other diffusion models. In this section, we use the Omni-ID encoder with IP-Adapter on FLUX Schnell [4] and SD15 [33] in the task of personalized text-to-image generation. Fig. III and Fig. IV demonstrates again the superiority of Omni-ID against other representations like CLIP and ArcFace.
Figure II: Qualitative comparisons to the state-of-the-art representations in controllable face generation. We compare Omni-ID with ArcFace [10] and CLIP [29] with 5 input images. To control each face in the grid, we drive the facial landmark of each identity by the same template. Our Omni-ID achieves superior identity preservation, captures nuanced details more faithfully, and demonstrates higher adaptivity to diverse poses and expressions.
Figure III: Qualitative comparisons with the state-of-the-art in personalized T2I generation using FLUX [4] as the base model. Our Omni-ID with IP-Adapter [47] without any other regularization (LoRA [21], ID loss [17], alignment loss [17]) achieves highest ID preservation. Different variants of IP-Adapter without LoRA are shown at the left side. The state-of-the-art PuLID-FLUX-v0.9.1 achieves lower face quality compared to Omni-ID. Omni-ID also works well on FLUX Schnell model, which generates each sample by $4$ denoising steps.
Figure IV: Qualitative comparisons to the state-of-the-art in personalized T2I generation using Stable Diffusion [33] as the base model. IPA-Full, IPA-Plus, and our IPA-Omni-ID use SD15 [33] as the base model, generating $512\times 512$ resolution samples. InstantID [43] and PhotoMakerV2 [25] use SDXL [28] as the base model, generating $1024\times 1024$ samples, which are resized to $512\times 512$ to show with other methods side by side. Our Omni-ID with IP-Adapter without any other regularization achieves the highest ID preservation.